OpenStudy (anonymous):

Best answer? Will give medal. I've been looking for an explanation to this question for a while. Anybody know how to get the answer? Solve the compound inequality. Write solution as a compound inequality. -4/3 is less than or equal to 1/7 w-3/4<1

OpenStudy (anonymous):

-4/3 <= 1/7w - 3/4 < 1 We want to isolate that w in the middle, right? So first get rid of that -3/4 by adding 3/4 to the left, middle, and right sides of the inequality. What do you get?

OpenStudy (anonymous):

-7/12 is less than or equal to 25/28w<1 3/4?

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

No. Your middle is a little messed up. And write the right side as a single fraction -4/3 + 3/4 <= 1/7w - 3/4 + 3/4 < 1 + 3/4 -7/12 <= 1/7w < 7/4

OpenStudy (anonymous):

Now get rid of the 1/7 by multiplying all 3 sides by 7

OpenStudy (anonymous):

-4 1/2 <= 1w < 12 1/4?

OpenStudy (anonymous):

Not sure what you meant about writing the right side a single fraction....